# Hawking Radiation Signatures from Primordial Black Holes Transiting the Inner Solar System: Prospects for Detection

### v1

@@ I Introduction
Primordial black holes (PBH) are theorized to form from the collapse of primordial density perturbations in the early universe. Those that form with sufficient initial mass, $M_{i}\gtrsim 5.34\times 10^{14}\,{\rm g}$, would persist to the present day and contribute to the dark matter (DM) density $\Omega_{\rm DM}$ Zel’dovich and Novikov (1966); Hawking (1971); Carr and Hawking (1974); Carr (1976); Khlopov et al. (1985); Khlopov (2010); Escrivà et al. (2024). Various cosmological models predict the existence of PBHs over a wide present-day mass range which, in principle, could span approximately $40$ orders of magnitude, from $10^{14}\,{\rm g}$ to $10^{55}\,{\rm g}$ Carr et al. (2010); Carr and Kuhnel (2020); Carr et al. (2021); Carr and Kuhnel (2022); Escrivà et al. (2024); Özsoy and Tasinato (2023). This translates to $10^{-19}M_{\odot}\lesssim M\lesssim 10^{22}M_{\odot}$. Despite many experimental constraints on the PBH dark matter fraction $f_{\rm PBH}$, there remains an open “asteroid-mass” window within which PBHs of mass $10^{17}\,{\rm g}\lesssim M\lesssim 10^{23}\,{\rm g}$ may comprise up to 100% of the cold dark matter (CDM) Carr et al. (2010); Carr and Kuhnel (2020); Carr et al. (2021); Carr and Kuhnel (2022); Green and Kavanagh (2021); Carr et al. (2024); Gorton and Green (2024); Escrivà et al. (2024); Özsoy and Tasinato (2023). We are motivated to experimentally search for PBHs not only because they could make up a significant fraction of the dark matter, but because the presence or absence of PBHs in our universe, even for $f_{\rm PBH}\ll 1$, can allow us to constrain cosmological models and to probe fundamental and beyond-Standard-Model (BSM) physics.
@@ I.1 Motivation
Primordial black holes garnered new attention as a dark matter candidate over the last decade in the wake of LIGO-Virgo-KAGRA (LVK) collaboration data and James Webb Space Telescope (JWST) observations. Some physicists Clesse and Garcia-Bellido (2022); Huang et al. (2024) have interpreted gravitational-wave signatures from binary black hole mergers with candidates in the lower Abbott et al. (2023a) and pair-instability Abbott et al. (2020) mass gaps as potential PBH candidates. Additionally, several analyses of public LVK data claim to observe sub-solar mass black holes Phukon et al. (2021); Morras et al. (2023), which would be excellent PBH candidates. The LVK collaboration, however, reports no such candidates with sufficient significance in their second and third observing runs Abbott et al. (2023b, 2018). Furthermore, JWST observations of galaxies Castellano et al. (2022, 2023); Adams et al. (2023); Atek et al. (2023); Harikane et al. (2023); Naidu et al. (2022) and candidate active galactic nuclei (AGN) Bogdan et al. (2024); Natarajan et al. (2024); Kovacs et al. (2024); Larson et al. (2023); Maiolino et al. (2024) with redshift $z\gtrsim 10$ are in tension with the $\Lambda$CDM standard model of cosmology Menci et al. (2022). These observations have reinvigorated interest in PBHs as seeds for early galaxies and the supermassive black holes at the centers of their AGNs Dolgov (2024); Hütsi et al. (2023); Silk et al. (2024); Liu and Bromm (2022).
The $\Lambda$CDM model predicts that about $26\%$ of the energy density of the universe is comprised of non-relativistic, non-baryonic matter with extremely weak coupling to Standard Model particles Aghanim et al. (2020). Whereas the evidence for the existence and relative fraction of DM is very strong, as yet there does not exist any decisive evidence to support various DM particle candidates, such as axions, axion-like particles, weakly interacting massive particles (WIMPs), or sterile neutrinos Boveia and Doglioni (2018); Weltman et al. (2020); Agrawal et al. (2021); Cooley et al. (2022); Workman et al. (2022). Primordial black holes also satisfy the $\Lambda$CDM dark matter criteria, and their formation and properties can be explained entirely within current theories of general relativity, $\Lambda\text{CDM}$, inflation, and the Standard Model. Most models predict that PBHs formed before baryogenesis and thus contribute to the non-baryonic dark matter density $\Omega_{\text{DM}}$ rather than to the baryonic density $\Omega_{b}$—which can be at most $5\%$ of the critical energy density Carr and Kuhnel (2020). Because PBHs with masses within the asteroid-mass range (or larger) are expected to be uncharged under any Standard Model gauge group in the present-day epoch Carter (1974); Alonso-Monsalve and Kaiser (2024), they can only interact gravitationally with SM particles. Furthermore, due to their large masses, PBHs are inherently non-relativistic—“cold”—objects capable of coalescing to form large-scale structures in the universe.
Primordial black holes also provide a useful tool with which to investigate cosmological history. Proposed PBH formation mechanisms include the collapse of primordial over-densities, pressure reduction in the QCD epoch, cosmic phase transitions and domain-wall collapse, self-intersection of cosmic strings, bubble collisions, and collapse of a scalar field. (See Refs. Carr and Kuhnel (2022); Green and Kavanagh (2021); Escrivà et al. (2024); Özsoy and Tasinato (2023) for a thorough discussion of possible PBH formation processes.) The parameters of a specific cosmological model determine the shape of the present-day PBH number density distribution, $dn/dM$, and the fraction of DM comprised of PBHs, $f_{\rm PBH}=\Omega_{\rm PBH,0}/\Omega_{\rm DM,0}$. Working backward from experimental constraints on $f_{\rm PBH}$, we can constrain cosmological models and probe the physics of the early universe and the shape of the primordial power spectrum.
@@ I.2 Evaporation Constraints
Across their entire mass range, PBHs should interact with known matter in predictable ways. These interactions manifest as observational signatures, which can be used to directly detect PBHs or to place constraints on $f_{\rm PBH}$. Potential signatures include Hawking emission, accretion, dynamical effects, large-scale structure formation, gravitational lensing, and gravitational waves. See Refs. Escrivà et al. (2024); Carr et al. (2010); Carr and Kuhnel (2020); Green and Kavanagh (2021); Carr et al. (2021); Carr and Kuhnel (2022); Özsoy and Tasinato (2023); Carr et al. (2024); Gorton and Green (2024) for a thorough review of current constraints from these signatures and others. We are interested in developing a new PBH detection signature and the requisite data analysis techniques to probe the open asteroid-mass window, $10^{17}\text{ g}\leq M\leq 10^{23}$ g. Reaching this notoriously inaccessible PBH mass range requires the development of novel signatures because these black holes have Schwarzschild radii, $R_{s}\equiv 2GM/c^{2}$, smaller than a micron, thus rendering known gravitational lensing techniques to be much more difficult to apply.
@@
The black holes we discuss in this article have initial formation-time masses $5.34\times 10^{14}\text{ g}\leq M_{i}\leq 5\times 10^{17}$ g, a window we refer to as the “$e^{\pm}$-production mass range,” which corresponds to radii between approximately $1-1000\,{\rm fm}$. We are focusing our initial efforts on positron Hawking emission, and therefore on PBHs with present-day masses $M\leq 5\times 10^{17}\,{\rm g}$, because positrons have low astrophysical backgrounds and the AMS-02 experiment has a large acceptance and provides over a decade of high-precision time-series positron flux data.
@@
Hawking radiation provides useful evaporation constraints on the dark matter fraction of low-mass PBHs with $M\lesssim 10^{17}{\rm g}$. Hawking’s semi-classical formulation of black hole evaporation Hawking (1975, 1974) predicts that a PBH should emit all SM particles during the course of its lifetime. PBHs that form within the $e^{\pm}$-producing range, $5.34\times 10^{14}\leq M_{i}\leq 5\times 10^{17}$ g, will emit photons, neutrinos, light charged leptons, and pions peaking in the MeV-GeV energy range, while PBHs in the asteroid-mass window will emit photons and neutrinos peaking in the eV-MeV range. PBHs at the very end of their lives will rapidly lose their remaining mass and effectively explode by emitting high fluxes of all 17 Standard Model particles—including heavy hadrons, leptons, and massive bosons. For example, as discussed in section II.2.2, a PBH with mass $M_{*}\equiv 5.34\times 10^{14}$ g has a lifetime equal to the current age of the universe, while a PBH with mass $6\times 10^{10}$ g has a lifetime of approximately 1 day and a temperature $T\approx 176\,{\rm GeV}$, which is hot enough to emit top quarks and Higgs bosons. PBH explosions could also be sources of ultra-high-energy cosmic rays Boluna et al. (2024); Baker et al. (2025a); Klipfel and Kaiser (2025); Baker et al. (2025b); Airoldi et al. (2025) and heavy antimatter cosmic rays Fisher et al. (2025).
@@ II.1 Hawking Radiation
Hawking radiation Hawking (1974, 1975); Page (1976a, b, 1977) is not limited to the emission of massless particles. Provided that a black hole is small enough and therefore hot enough, it can emit all 17 Standard Model particles as fundamental degrees of freedom. These fundamental emission rates are referred to as primary emission spectra. To determine the feasibility of detecting Hawking radiation from primordial black holes transiting past Earth with space-based cosmic ray experiments like AMS, we must accurately simulate the PBH emission spectra for photons, electrons, and positrons. Because our technique probes hot, low-mass PBHs, it is critical that we account for QCD jets, hadronization, and particle decays associated with the primary emission of quarks, muons and other short-lived particles MacGibbon and Webber (1990). Jet fragmentation and particle decays will contribute to the net emission rates—referred to as secondary emission spectra—of stable, detectable fundamental and composite particles. We therefore will use secondary rather than primary emission spectra throughout our analysis.
@@ II.1.1 Primary Emission
The important physics enters through the greybody factor, $\Gamma_{s}$, a dimensionless quantity which is generally a function of the black hole parameters $\Omega,\ \Phi,\ M,$ and $\kappa$ and the particle parameters $Q,\ m,$ and $s$. The greybody factor is defined in terms of the absorption cross section $\sigma_{s}$ associated with scattering a Standard Model field of mass $m$ and spin $s$ off a black hole of mass $M$ in curved spacetime MacGibbon and Webber (1990):
@@ II.2.1 PBH Mass Functions
[…]
@@
Equation (9) implies that PBHs can form with arbitrarily small masses, if the (averaged) compaction is close to the critical threshold. Meanwhile, causality restricts PBHs to form with masses no larger than the Hubble mass at the time of formation. In fact, the typical mass at the peak of the distribution at the time of formation satisfies $\bar{M}(t_{i})\simeq 0.2\,M_{H}(t_{i})$ Carr (1975); Niemeyer and Jedamzik (1998); Green and Liddle (1999); Kühnel et al. (2016). These features of the critical collapse process generically yield PBH mass distributions that have a power-law tail for masses $M_{i}<\bar{M}(t_{i})$ and an exponential cut-off for masses $M_{i}\gtrsim\bar{M}(t_{i})$.
@@ II.2.2 PBH Mass Loss
[…]
@@
[…]
@@
It is common to make the further assumption that $f(M)=f(M_{i})$, which allows one to solve directly for $M(t)$ Mosbech and Picker (2022):
@@ II.2.3 Present-Day Mass Functions
[…]
@@
[…]
@@
To derive a general expression for $\psi(M,t)$ at some time $t>t_{i}$, we take advantage of the conservation of the total number of PBHs within a certain mass bin in the present day. The number $N$ of PBHs with mass in some range $[M_{1},\,M_{2}],\,\,M_{1}>0,$ at time $t$ is given by the integral over the function $\phi(M)$:
[…]
@@
$\begin{split}\phi(M,t)&\approx M^{2}\phi\left(M_{i}(M,t),t_{i}\right)\\ &\,\,\,\,\,\,\times\left(M^{3}+3Atf(M_{i}(M,t))\right)^{-2/3}.\\ \end{split}$ (29)
@@
$\begin{split}\psi(M,t)&\approx\frac{M^{3}}{M_{i}(M,t)}\psi\left(M_{i}(M,t),t_{i}\right)\\ &\,\,\,\,\,\times\left(M^{3}+3Atf(M_{i}(M,t))\right)^{-2/3}.\\ \end{split}$ (31)

### v2

@@ I Introduction
Primordial black holes (PBH) are theorized to form from the collapse of primordial density perturbations in the early universe. Those that form with sufficient initial mass, $M_{i}\gtrsim 5.364\times 10^{14}\,{\rm g}$, would persist to the present day and contribute to the dark matter (DM) density $\Omega_{\rm DM}$ Zel’dovich and Novikov (1966); Hawking (1971); Carr and Hawking (1974); Carr (1976); Khlopov et al. (1985); Khlopov (2010); Escrivà et al. (2024). Various cosmological models predict the existence of PBHs over a wide present-day mass range which, in principle, could span approximately $40$ orders of magnitude, from $10^{14}\,{\rm g}$ to $10^{55}\,{\rm g}$ Carr et al. (2010); Carr and Kuhnel (2020); Carr et al. (2021); Carr and Kuhnel (2022); Escrivà et al. (2024); Özsoy and Tasinato (2023). This translates to $10^{-19}M_{\odot}\lesssim M\lesssim 10^{22}M_{\odot}$. Despite many experimental constraints on the PBH dark matter fraction $f_{\rm PBH}$, there remains an open “asteroid-mass” window within which PBHs of mass $10^{17}\,{\rm g}\lesssim M\lesssim 10^{23}\,{\rm g}$ may comprise up to 100% of the cold dark matter (CDM) Carr et al. (2010); Carr and Kuhnel (2020); Carr et al. (2021); Carr and Kuhnel (2022); Green and Kavanagh (2021); Carr et al. (2024); Gorton and Green (2024); Escrivà et al. (2024); Özsoy and Tasinato (2023); Boybeyi et al. (2025). We are motivated to experimentally search for PBHs not only because they could make up a significant fraction of the dark matter, but because the presence or absence of PBHs in our universe, even for $f_{\rm PBH}\ll 1$, can allow us to constrain cosmological models and to probe fundamental and beyond-Standard-Model (BSM) physics.
@@ I.1 Motivation
[…]
Primordial black holes also provide a useful tool with which to investigate cosmological history. PBHs would form from the collapse of primordial overdensities. Several mechanisms have been proposed that could amplify primordial overdensities on appropriate length-scales to seed PBH formation, including a sudden change in the effective sound speed at the time of the QCD phase transition, domain-wall collapse, self-intersection of cosmic strings, or bubble collisions from strongly first-order phase transitions Jedamzik and Niemeyer (1999); Liu et al. (2022); Baker et al. (2025a, b); Lewicki et al. (2023); Gonçalves et al. (2025). (See Refs. Carr and Kuhnel (2022); Green and Kavanagh (2021); Escrivà et al. (2024); Özsoy and Tasinato (2023) for a thorough discussion of possible PBH formation processes.)
Among the best-studied formation mechanisms involve density perturbations that were amplified during a phase of early-universe inflation. When overdensities above some critical threshold re-enter the Hubble radius following the end of inflation, they induce gravitational collapse into black holes. In these scenarios, the overdensities are associated with a peak in the spectrum of primordial gauge-invariant curvature perturbations Niemeyer and Jedamzik (1998a); Green and Liddle (1999); Young et al. (2019); Escrivà et al. (2020); Gow et al. (2021); Musco et al. (2021). (See especially the reviews in Refs. Gundlach and Martin-Garcia (2007); Escrivà (2022).) Such peaks can form from dynamics during inflation that depart from ordinary slow-roll evolution, such as a transient phase of ultra-slow-roll dynamics Kinney (2005); Garcia-Bellido and Morales (2017); Motohashi and Hu (2017); Di and Gong (2018); Pattison et al. (2021); Geller et al. (2022) or brief tachyonic growth of perturbations on short length-scales associated with rapid turns in field space (within multifield models) Garcia-Bellido et al. (1996); Lyth (2011); Halpern et al. (2015); Clesse and García-Bellido (2015); Fumagalli et al. (2023); Braglia et al. (2020); Palma et al. (2020); Lorenzoni et al. (2025). Recent works have demonstrated that such dynamics can arise within models of the early universe that incorporate realistic features from high-energy physics, including multiple interacting fields, with less fine-tuning of parameters than typically required in simpler single-field constructions Qin et al. (2023); Lorenzoni et al. (2025).
The parameters of a specific cosmological model determine the shape of the present-day PBH number density distribution, $dn/dM$, and the fraction of DM comprised of PBHs, $f_{\rm PBH}=\Omega_{\rm PBH,0}/\Omega_{\rm DM,0}$. Working backward from experimental constraints on $f_{\rm PBH}$, we can constrain cosmological models and probe the physics of the early universe and the shape of the primordial power spectrum.
@@ I.2 Evaporation Constraints
Across their entire mass range, PBHs should interact with known matter in predictable ways. These interactions manifest as observational signatures, which can be used to directly detect PBHs or to place constraints on $f_{\rm PBH}$. Potential signatures include Hawking emission, accretion, dynamical effects, large-scale structure formation, gravitational lensing, and gravitational waves. See Refs. Escrivà et al. (2024); Carr et al. (2010); Carr and Kuhnel (2020); Green and Kavanagh (2021); Carr et al. (2021); Carr and Kuhnel (2022); Özsoy and Tasinato (2023); Carr et al. (2024); Gorton and Green (2024); Boybeyi et al. (2025) for a thorough review of current constraints from these signatures and others. We are interested in developing a new PBH detection signature and the requisite data analysis techniques to probe the open asteroid-mass window, $10^{17}\text{ g}\leq M\leq 10^{23}$ g. Reaching this notoriously inaccessible PBH mass range requires the development of novel signatures because these black holes have Schwarzschild radii, $R_{s}\equiv 2GM/c^{2}$, smaller than a micron, thus rendering known gravitational lensing techniques to be much more difficult to apply.
@@
The black holes we discuss in this article have initial formation-time masses $5.364\times 10^{14}\text{ g}\leq M_{i}\leq 5\times 10^{17}$ g, a window we refer to as the $e^{\pm}$-production mass range,” which corresponds to radii between approximately $1-1000\,{\rm fm}$. We are focusing our initial efforts on positron Hawking emission, and therefore on PBHs with present-day masses $M\leq 5\times 10^{17}\,{\rm g}$, because positrons have low astrophysical backgrounds and the AMS-02 experiment has a large acceptance and provides over a decade of high-precision time-series positron flux data.
@@
Hawking radiation provides useful evaporation constraints on the dark matter fraction of low-mass PBHs with $M\lesssim 10^{17}{\rm g}$. Hawking’s semi-classical formulation of black hole evaporation Hawking (1975, 1974) predicts that a PBH should emit all SM particles during the course of its lifetime. PBHs that form within the $e^{\pm}$-producing range, $5.364\times 10^{14}\leq M_{i}\leq 5\times 10^{17}$ g, will emit photons, neutrinos Klipfel and Kaiser (2025); Lysyy et al. (2024), light charged leptons, and pions peaking in the MeV-GeV energy range, while PBHs in the asteroid-mass window will emit photons and neutrinos peaking in the eV-MeV range. PBHs at the very end of their lives will rapidly lose their remaining mass and effectively explode by emitting high fluxes of all 17 Standard Model particles—including heavy hadrons, leptons, and massive bosons. For example, as discussed in section II.2.2, a PBH with mass $M_{*}\equiv 5.364\times 10^{14}$ g has a lifetime equal to the current age of the universe, while a PBH with mass $6\times 10^{10}$ g has a lifetime of approximately 1 day and a temperature $T\approx 176\,{\rm GeV}$, which is hot enough to emit top quarks and Higgs bosons. PBH explosions could also be sources of ultra-high-energy cosmic rays Boluna et al. (2024); Baker et al. (2025c); Klipfel and Kaiser (2025); Baker et al. (2025b); Airoldi et al. (2025) and heavy antimatter cosmic rays Fisher et al. (2025).
@@ II.1 Hawking Radiation
A black hole can couple to every form of matter because of the universal nature of gravitation. Therefore, Hawking radiation Hawking (1974, 1975); Page (1976a, b, 1977) is not limited to the emission of photons. Provided that a black hole is small enough and therefore hot enough, it can emit all 17 Standard Model particles as fundamental degrees of freedom. These fundamental emission rates are referred to as primary emission spectra. To determine the feasibility of detecting Hawking radiation from primordial black holes transiting past Earth with space-based cosmic ray experiments like AMS, we must accurately simulate the PBH emission spectra for photons, electrons, and positrons. Because our technique probes hot, low-mass PBHs, it is critical that we account for QCD jets, hadronization, and particle decays associated with the primary emission of quarks, muons and other short-lived particles MacGibbon and Webber (1990). Jet fragmentation and particle decays will contribute to the net emission rates—referred to as secondary emission spectra—of stable, detectable fundamental and composite particles. We therefore will use secondary rather than primary emission spectra throughout our analysis.
@@ II.1.1 Primary Emission
The important physics enters through the greybody factor, $\Gamma_{s}$, a dimensionless quantity that governs the interaction between the black hole and the emitted particles, which is generally a function of the black hole parameters $\Omega,\ \Phi,\ M,$ and $\kappa$ and the particle parameters Q,m,Q,\ m, and $s$. The greybody factor is defined in terms of the absorption cross section $\sigma_{s}$ associated with scattering a Standard Model field of mass $m$ and spin $s$ off a black hole of mass $M$ in curved spacetime MacGibbon and Webber (1990):
@@ II.2.1 PBH Mass Functions
where $\psi(M_{i})$ satisfies $\int dM_{i}\,\psi(M_{i})=1$ and determines the shape of the initial distribution. Here $n_{{\rm PBH},i}$ is the comoving PBH number density at formation time.
@@
Equation (9) implies that PBHs can form with arbitrarily small masses, if the (averaged) compaction is close to the critical threshold. Meanwhile, causality restricts PBHs to form with masses no larger than the Hubble mass at the time of formation. In fact, the typical mass at the peak of the distribution at the time of formation satisfies $\bar{M}(t_{i})\simeq 0.2\,M_{H}(t_{i})$ Carr (1975); Niemeyer and Jedamzik (1998b); Green and Liddle (1999); Kühnel et al. (2016). These features of the critical collapse process generically yield PBH mass distributions that have a power-law tail for masses $M_{i}<\bar{M}(t_{i})$ and an exponential cut-off for masses $M_{i}\gtrsim\bar{M}(t_{i})$.
@@ II.2.2 PBH Mass Loss
$\begin{split}\frac{dM}{dt}&=-\sum_{j}g_{j}\int_{m_{\rm eff},j}^{\infty}dQ\,\frac{Q}{c^{2}}\frac{d^{2}N^{(1)}_{j}}{dtdQ}\\ &=-A\,\frac{f(M)}{M^{2}},\end{split}$ (14)
@@
$\begin{split}\int\frac{M^{2}}{f(M)}dM&=\frac{1}{3}\frac{M^{3}}{f(M)}-\int\frac{1}{3}M^{3}\frac{d}{dM}f(M)^{-1}dM\\ &=\frac{1}{3}\frac{M^{3}}{f(M)}+\int\frac{1}{3}\frac{M^{3}}{f(M)^{2}}\frac{df(M)}{dM}dM\\ &\approx\frac{1}{3}\frac{M^{3}}{f(M)},\end{split}$ (17)
@@
(See also Ref. Cai et al. (2023).) It is common to make the further assumption that $f(M)=f(M_{i})$, which allows one to solve directly for $M(t)$ Mosbech and Picker (2022):
@@ II.2.3 Present-Day Mass Functions
The simplest way to construct an approximate present-day PBH mass function is to truncate the formation-time mass function $\psi(M_{i})$ for $M_{i}\leq M_{*}$. Assuming that all PBHs with $M_{i}<M_{*}$ have evaporated, one could then define a present-day mass function $\psi(M,t_{0})$ via $\psi(M,t_{0})\propto\psi(M_{i})\,\Theta(M_{i}-M_{*}),$ where $\Theta(x)$ is the Heaviside step function. Such a distribution could be normalized by requiring $\int_{M_{*}}^{\infty}dM\,\psi(M,t_{0})=1$. In practice, for PBH mass functions that peak at $\bar{M}\gg M_{*}$, one may neglect the evolution of the PBH population over time and simply set $\psi(M_{i})\simeq\psi(M)$, which is commonly done in the literature.
@@
This function is defined such that $\phi(M_{i})dM_{i}$ is the fraction of PBHs with mass $M_{i}$ at formation time, and satisfies $\int dM_{i}\,\phi(M_{i})=1$.
@@
To derive a general expression for $\psi(M,t)$ at some time $t>t_{i}$, we take advantage of the conservation of the total number of PBHs within a certain mass bin in the present day. The number $N$ of PBHs with mass in some range $M_{1},\,M_{2}],\,\,M_{1}>0,$ at time $t$ is given by the integral over the function $\phi(M)$:
$\begin{split}N&=n_{{\rm PBH},i}\int_{M_{1}}^{M_{2}}dM\,\phi(M,t)\\ &=n_{{\rm PBH},i}\int_{M_{i}(M_{1},t)}^{M_{i}(M_{2},t)}dM\,\phi(M(M_{i},t),t_{i})\left[\frac{dM}{dM_{i}}\right]^{-1},\end{split}$ (27)
@@
[…]
@@
[…]